# Harder-Narasimhan filtrations of decorated vector bundles

### v1

@@ Abstract.
A decorated vector bundle is a vector bundle equipped with a reduction of structure group to a complex reductive subgroup $G\subseteq\mathbf{GL}(r;\mathbb{C})$. Examples include symplectic and special-orthogonal vector bundles, as well as vector bundles with trivial determinants. In this expository paper, we provide direct constructions of Harder-Narasimhan filtrations of symplectic and special-orthogonal vector bundles, and use them to construct canonical reductions as done by Atiyah and Bott. We compare these canonical reductions to those constructed by Biswas and Holla. Lastly, we set up the obstruction theory necessary to define Harder-Narasimhan types of principal bundles, and stratify the moduli stack of principal $G$-bundles.
@@ 1. Semistable bundles
We compare the slope-semistability of vector bundles from Mumford in [mumf_projectiveinvar] with the semistability of principal bundles from Ramanathan in [ramanathan_stableprincipalbundles]. Following Hyeon-Murphy in [hyeon-murphy_noteonthestabilityofprincbund], we recall that they coincide when the structure group is $\mathbf{GL}(r,\mathbb{C})$.
@@ 1.1. Slope-semistability
In [mumf_projectiveinvar], Mumford introduced slope-semistability and slope-stability to study moduli problems of vector bundles, using geometric invariant theory.
@@
To generalize stability to principal-$G$-bundles, where $G$ is a complex reductive group, we wish to rephrase it in terms of principal-$\mathbf{GL}(r,\mathbb{C})$-bundles. As $\mathbf{GL}(r,\mathbb{C})$ is a reductive group, we consider the Cartan subgroup $T$ of $\mathbf{GL}(r,\mathbb{C})$ of diagonal matrices, and the Borel subgroup $B$ of upper-right triangular matrices, such that $T\subseteq B\subseteq\mathbf{GL}(r,\mathbb{C})$. We call those the standard Cartan and Borel subgroups of $\mathbf{GL}(r,\mathbb{C})$. A standard parabolic subgroup $P\subseteq\mathbf{GL}(r,\mathbb{C})$ is a subgroup that contains $B$. Such subgroups consist of upper-right block matrices and correspond uniquely to subsets $I$ of the set of simple roots $\triangle$ induced by $B$:
@@
where $A=(A_{ij})_{1\leqslant i,j\leqslant r}$ for every $r\times r$ matrix with coefficients in $\mathbb{C}$. We henceforth index standard parabolics as $P_{I}$ and we refer to [borel_linearalgebraicgroups] for a review of Cartan, Borel, and parabolic subgroups of reductive groups.
@@
For each subbundle $F$, the corresponding section $s:X\rightarrow\mathrm{Fr}(E)/P_{I}$ gives frames of $E$ that extend frames of $F$. Given local trivializations of $E$, we can induce from $s$ a family of cocycles $(\sigma_{i,j})_{i,j\in J}$ of $E$, compatible with $F$ as in Remark 1.2. Through Lemma 1.3, the natural way to generalize slope-stability to principal-$G$-bundles is by imposing slope conditions on reductions to standard maximal parabolics $P$ of $G$. For a decorated vector bundle $(E,s)$, i.e., a vector bundle equipped with a reduction $s:X\rightarrow\mathrm{Fr}(E)/G$ of structure group $G\subseteq\mathbf{GL}(r,\mathbb{C})$, it means that the slope conditions from Definition 1.1 are tested only on certain subbundles $F\subsetneq E$ that respect the decoration. We first work out the $G=\mathbf{SL}(r,\mathbb{C})$ case, corresponding to vector bundles with trivial determinants, i.e., special vector bundles.
@@
For a rank $r$ special vector bundle $(E,\tau)$, giving a section $\tau$ is equivalent to giving a reduction $s:X\rightarrow\mathrm{Fr}(E)/\mathbf{SL}(r,\mathbb{C})$ of the frame bundle. Hence, a special vector bundle is a decorated vector bundle. Just as vector bundles $E$, of rank $r$, correspond to principal-$\mathbf{GL}(r,\mathbb{C})$-bundles through their frame bundles $\mathrm{Fr}(E)$, special vector bundles $(E,\tau)$, of rank $r$, correspond to principal-$\mathbf{SL}(r,\mathbb{C})$-bundles through their special frame bundles $\mathrm{Fr}_{\mathbf{SL}}(E,\tau)$.
@@
For a special vector bundle $(E,\tau)$ of rank $r$, the special frame bundle $\mathrm{Fr}_{\mathbf{SL}}(E,\tau)$ is a principal-$\mathbf{SL}(r,\mathbb{C})$-bundle, whose fiber at $x\in X$ is given by:
@@
The right-$\mathbf{SL}(r,\mathbb{C})$-action on the fiber of $x\in X$ is given by:
@@
Following the discussion after Lemma 1.3, stability for vector bundles $E$ and special vector bundles $(E,\tau)$ should be equivalent since Lemmas 1.3 and 1.6 give the analogous correspondences. The decoration of a reduction of $E$ to $\mathbf{SL}(r,\mathbb{C})$ does not restrict the set of subbundles $F\subsetneq E$ with respect to which one tests the slope conditions. This means that the stability and semistability conditions for decorated vector bundles $(E,\tau)$, where $\tau$ is a non-vanishing section of $\det(E)$, are exactly that of stability and semistability of the underlying vector bundle $E$. Observe that $\mathbf{SL}(r,\mathbb{C})$ is the commutator subgroup $\mathbf{GL}(r,\mathbb{C}),\mathbf{GL}(r,\mathbb{C})]$ of $\mathbf{GL}(r,\mathbb{C})$. In Lemma 1.13, we will see the same phenomenon as Lemma 1.6, that the stability of principal-$[G,G]$-bundles and the stability of their extensions to principal-$G$-bundles are equivalent. Note that the commutator subgroup $[G,G]$ is always semisimple and has the same root system as $G$.
@@ 1.2. Ramanathan-semistability
We review Ramanathan’s work in [ramanathan_stableprincipalbundles], including his definition of semistability and stability for principal-$G$-bundles, where $G$ is complex and reductive. Let $\xi$ be a holomorphic principal-$G$-bundle on a compact Riemann surface $X$. For a reduction $s:X\rightarrow\xi/H$ of $\xi$ to a subgroup $H\subseteq G$, we can pullback the bundle $\xi\rightarrow\xi/H$ by $s$ to obtain a principal-$H$-bundle $s^{*}\xi$ on $X$. We also refer to $s^{*}\xi$ as the reduction of $\xi$ to $H$. Following [ramanathan_stableprincipalbundles], we define the vertical tangent bundle $V_{\xi/H}$, which is the vector bundle on $\xi/H$ consisting of tangent vectors along the fibers of $\xi/H\rightarrow X$.
@@
Let $\xi$ be a holomorphic principal-$G$-bundle on a compact Riemann surface $X$.
@@
as seen in [ramanathan_stableprincipalbundles, Remark 2.2]. Since $G$ is reductive, there exists a nondegenerate adjoint-invariant form on $\mathrm{Lie}(G)$ such that $\mathrm{ad}(\xi)$ is self-dual, so $\mathrm{ad}(\xi)$ has degree 0. From the short exact sequence, we have $\deg(s^{*}V_{\xi/P})=-\deg(\mathrm{ad}(s^{*}\xi))$. This can be used as an alternate characterization of Ramanathan-(semi)-stability. Namely, $\xi$ is Ramanathan-(semi)-stable when for all reductions $s:X\rightarrow\xi/P$ to maximal parabolic subgroups $P$ of $G$, we have $\deg(\mathrm{ad}(s^{*}\xi))(\leqslant)0$.
@@
Once we fix a Borel subgroup $B$, to verify Ramanathan-(semi)-stability, it suffices to verify the inequality for reductions to maximal standard parabolic subgroups of $G$. For a principal-$G$-bundle $\xi$ with this property and a reduction $s:X\rightarrow\xi/P$ to any maximal parabolic subgroup $P\subsetneq G$, we have that:
@@
    There exists $g\in G$, such that $gPg^{-1}=P_{I}$ for some standard maximal parabolic subgroup $P_{I}\subsetneq G$. See [*][IV. 11.1 Theorem, IV. 11.3 Corollary]borel_linearalgebraicgroups.
@@
The next lemma is from [*][Proposition 1]hyeon-murphy_noteonthestabilityofprincbund, which we use to recall the equivalence between slope-(semi)-stability and Ramanathan-(semi)-stability. Let $E$ be a vector bundle of rank $r$.
@@
We can now prove the following theorem, from [*][Corollary 1]hyeon-murphy_noteonthestabilityofprincbund.
@@
We can prove Theorem 1.12 in the same way as Theorem 1.11. Alternatively, we can invoke the following lemma, which generalizes the connection between Theorems 1.11 and 1.12 and shows why the notion of stability and semistability of a principal $[G,G]$-bundle coincides with that of the asssociated $G$-bundle.
@@
For a principal-$[G,G]$-bundle $\xi_{1}$, the following are equivalent:
@@
Let $P_{1}$ be a maximal parabolic subgroup of $[G,G]$. Let $P=P_{1}\mathbf{R}(G)$ be the subgroup of $G$ generated by $P_{1}$ and the radical $\mathbf{R}(G)$ of $G$, as defined in [borel_linearalgebraicgroups, IV. 11.21], then $P$ is a maximal parabolic subgroup of $G$, since $G=[G,G]\mathbf{R}(G)$ following [borel_linearalgebraicgroups, IV. 14.2 Proposition], and the fact that $\mathbf{R}(G)$ is the connected component of the center $\mathbf{Z}(G)$ from [borel_linearalgebraicgroups, IV. 11.21 Proposition]. Furthermore, we have a $[G,G]$-equivariant isomorphism of complete varieties:
@@
Lastly, we recall an important use of Remark 1.8. The adjoint bundle of a principal-$G$-bundle determines its Ramanathan-semistability. To prove this, we use [ramanan-ramanathan_someremarksoninstabilityflag, Theorem 3.18].
@@
Let $\varphi:G\rightarrow H$ be a homomorphism of complex reductive groups that maps the radical $\mathbf{R}(G)$ into $\mathbf{R}(H)$. For a Ramanathan-semistable principal-$G$-bundle $\xi$, the extension $\xi(H)$ of $\xi$ to $H$ is Ramanathan-semistable.
This theorem, proven in [ramanan-ramanathan_someremarksoninstabilityflag] by defining admissible reductions, will be helpful to us in later sections as well. We now present the result of [ramanathan_moduliforprincipalbundlesI, 3.18 Corollary].
@@
A principal-$G$-bundle $\xi$ is Ramanathan-semistable if and only if its adjoint bundle $\mathrm{ad}(\xi)$ is slope-semistable.
@@
The forward implication is a consequence of Theorem 1.14: the radical of a reductive group is a subgroup of the center, from [borel_linearalgebraicgroups, IV. 11.21 Proposition], and therefore $\mathbf{R}(G)$ gets mapped to 0 under the adjoint representation, whose kernel is $\mathbf{Z}(G)$. Ramanathan-semistability is preserved in the frame bundle of the adjoint bundle $\mathrm{ad}(\xi)$, which is equivalent to slope-semistability due to Theorem 1.11.
@@
The reverse implication of Lemma 1.15 is also true for stability, as we can replace the inequalities with strict inequalities in the proof above. However, the same is not true for the forward implication. In [hyeon-murphy_noteonthestabilityofprincbund, Lemma 2], it is explained that for stability, the forward implication is already false if $G$ is not semisimple, as in the case of $G=\mathbf{GL}(r,\mathbb{C})$.

### v2

# arXiv:2511.17371v2[math.AG] 02 Mar 2026

@@ Abstract.
A decorated vector bundle is a vector bundle equipped with a reduction of structure group to a complex reductive subgroup $G\subseteq\mathbf{GL}(r,\mathbb{C})$. Examples include symplectic and special-orthogonal vector bundles, as well as vector bundles with trivial determinants. In this expository paper, we provide direct constructions of Harder-Narasimhan filtrations of symplectic and special-orthogonal vector bundles, and use them to construct canonical reductions in the sense of Atiyah and Bott. We compare these canonical reductions to those constructed by Biswas and Holla. Lastly, we set up the obstruction theory necessary to define Harder-Narasimhan types of principal bundles, and stratify the moduli stack of principal $G$-bundles.
1. Introduction
Geometric invariant theory (GIT) was developed by David Mumford as a way to construct quotients and solve moduli problems in algebraic geometry [Mum62]. The basic notion in this theory is that of a categorical quotient: given a $G$-action on a variety $X$, the goal is to find sufficient conditions on $G$ and $X$ so that an arbitrary $G$-invariant morphism $f:X\to Y$ factors through a canonical morphism $\pi:X\to X\sslash G$. Such a categorical quotient exists for instance when $G$ is reductive and $X$ is affine. But even in that case, the variety $X\sslash G$ is in general distinct from the orbit space $X/G$: to get a point in $X\sslash G$, one needs to further identify two orbits if their closures intersect, and this turns out to be equivalent to only keeping the closed orbits. More generally, to construct quotients of quasi-projective variety, one needs to impose a semistability condition on the orbits, which selects an open subset $X^{\mathrm{sst}}\subseteq X$, of which one can then construct a categorical quotient $X^{\mathrm{sst}}\sslash G$ [Ses72, Ses77]. Moreover, there is a dense open subset of $X^{\mathrm{sst}}\sslash G$ which is an orbit space: its points are in bijection with the so-called stable $G$-orbits in $X^{\mathrm{sst}}$.
As an application, one can construct moduli spaces of (semi)-stable vector bundles using GIT. Indeed, Mumford showed in [Mum62], that for vector bundles over a smooth projective curve, the GIT-(semi)-stability condition translates to a much more concrete slope-(semi)-stability condition (Definition 2.1). This was generalized by A. Ramanathan in [Ram75] to principal $G$-bundles for a reductive structure group $G$, where subbundles are replaced by parabolic reductions. When $G=\mathbf{GL}(r,\mathbb{C})$, Ramanathan recovers Mumford’s slope-(semi)-stability condition. Similarly, for vector bundles that are decorated with additional structure, such as a symplectic form or a non-degenerate quadratic form, it is possible to formulate Ramanathan’s semistability in terms of a slope condition, which is only tested on certain special subbundles (namely isotropic subbundles, in the case of symplectic and orthogonal vector bundles). For convenience, we provide an explicit proof of these facts in Theorems 3.6 and 3.13, building on the approach of Hyeon and Murphy in [HM04]. The point is that, for decorated vector bundles, stability and semistability can also be formulated using slope conditions that are tested only on the isotropic subbundles of these decorated vector bundles.
When studying vector bundles or principal bundles over an algebraic curve, a natural question is: How far is such a bundle from being semistable? For a vector bundle, the question is answered by looking at the Harder-Narasimhan filtration, first constructed in [HN75]. This filtration is uniquely determined by the condition that the successive quotients are semistable and have their slopes arranged in strictly decreasing order (see Theorem 4.1 for a proof of the Harder-Narasimhan theorem). For a principal bundle, the Harder-Narasimhan filtrations must be replaced by a reduction of structure group to a so-called canonical reduction to a uniquely determined parabolic subgroup. For complex reductive groups, canonical reductions were introduced by Atiyah and Bott in [AB83, $\S$10]. Over an arbitrary base field, the construction is due to K. Behrend in [Beh95] using complementary polyhedra (see also [BH04]). Since semistability of symplectic and orthogonal vector bundles can be characterized using slope-semistability, it is natural to try and characterize canonical reductions of such bundles in terms of certain special filtrations, and we write this out explicitly in Theorems 3.6 and 3.13. As we shall see, the parabolic subgroup associated to the canonical reduction of a principal $G$-bundle $\xi$ determines a filtration by isotropic sub-bundles of the decorated vector bundle associated to $\xi$ but, in order to relate this to the usual Harder-Narasimhan filtration of the underlying vector bundle, one needs to add the co-isotropic complements of these subbundles to the filtration. In our exposition, we follow the Atiyah-Bott approach and make systematic use of the special-orthogonal Harder-Narasimhan filtration of the adjoint bundle of a principal $G$-bundle (see Subsection 4.2). Then in Subsection 4.3, we compare this to the Biswas-Holla approach, in which the existence of a canonical reduction is proved using dominant characters. We unify the two approaches in Theorem 4.24. Important references for our exposition are the works of Friedman-Morgan [FM02] and Ho-Liu [HL09].
Since moduli problems can also be formulated in terms of moduli stacks, we can study the stratifications of the stack of principal bundles $Bun(G)$ induced by canonical reductions. This is related to the notion of $\Theta$-stratifications of moduli stacks introduced by D. Halpern-Leistner in [HLv5] and is currently a topic of active investigation in moduli theory (see for instance [AHLH23]). In Section 5, we show how canonical reductions induce strata of $Bun(G)$ labeled by Harder-Narasimhan types, and we make these types explicit for decorated vector bundles. When $G=\mathbf{GL}(r,\mathbb{C})$, the Harder-Narasimhan type of a vector bundle $E$ consists essentially of the topological invariants (rank and degree) of the successive quotients of the Harder-Narasimhan filtration of $E$.
@@ 2. Semistable bundles
We compare the slope-semistability of vector bundles from Mumford in [Mum62] with the semistability of principal bundles from Ramanathan in [Ram75]. Following Hyeon-Murphy in [HM04], we recall that they coincide when the structure group is $\mathbf{GL}(r,\mathbb{C})$.
@@ 2.1. Slope-semistability
In [Mum62], Mumford introduced slope-(semi)stability to study moduli problems of vector bundles, using geometric invariant theory.
@@
To generalize stability to principal $G$-bundles, where $G$ is a complex reductive group, we wish to rephrase it in terms of principal $\mathbf{GL}(r,\mathbb{C})$-bundles. As $\mathbf{GL}(r,\mathbb{C})$ is a reductive group, we consider the Cartan subgroup $T$ of $\mathbf{GL}(r,\mathbb{C})$ of diagonal matrices, and the Borel subgroup $B$ of upper-right triangular matrices, such that $T\subseteq B\subseteq\mathbf{GL}(r,\mathbb{C})$. We call those the standard Cartan and Borel subgroups of $\mathbf{GL}(r,\mathbb{C})$. A standard parabolic subgroup $P\subseteq\mathbf{GL}(r,\mathbb{C})$ is a subgroup that contains $B$. Such subgroups consist of upper-right block matrices and correspond uniquely to subsets $I$ of the set of simple roots $\triangle$ induced by $B$:
@@
where $A=(A_{ij})_{1\leqslant i,j\leqslant r}$ for every $r\times r$ matrix with coefficients in $\mathbb{C}$. We henceforth index standard parabolics as $P_{I}$, and refer to [Bor91] for a review of Cartan, Borel, and parabolic subgroups of reductive groups.
@@
For each subbundle $F$, the corresponding section $s:X\rightarrow\mathrm{Fr}(E)/P_{I}$ gives frames of $E$ that extend frames of $F$. Given local trivializations of $E$, we can induce from $s$ a family of cocycles $(\sigma_{i,j})_{i,j\in J}$ of $E$, compatible with $F$ as in Remark 2.2. Through Lemma 2.3, the natural way to generalize slope-stability to principal $G$-bundles is by imposing slope conditions on reductions to standard maximal parabolics $P$ of $G$. For a decorated vector bundle $(E,s)$, i.e., a vector bundle equipped with a reduction $s:X\rightarrow\mathrm{Fr}(E)/G$ of structure group $G\subseteq\mathbf{GL}(r,\mathbb{C})$, it means that the slope conditions from Definition 2.1 are tested only on certain subbundles $F\subsetneq E$ that respect the decoration. We first work out the $G=\mathbf{SL}(r,\mathbb{C})$ case, corresponding to vector bundles with trivial determinants, i.e., special vector bundles.
@@
For a rank $r$ special vector bundle $(E,\tau)$, giving a section $\tau$ is equivalent to giving a reduction $s:X\rightarrow\mathrm{Fr}(E)/\mathbf{SL}(r,\mathbb{C})$ of the frame bundle. Hence, a special vector bundle is a decorated vector bundle. Just as vector bundles $E$, of rank $r$, correspond to principal $\mathbf{GL}(r,\mathbb{C})$-bundles through their frame bundles $\mathrm{Fr}(E)$, special vector bundles $(E,\tau)$, of rank $r$, correspond to principal $\mathbf{SL}(r,\mathbb{C})$-bundles through their special frame bundles $\mathrm{Fr}_{\mathbf{SL}}(E,\tau)$.
@@
For a special vector bundle $(E,\tau)$ of rank $r$, the special frame bundle $\mathrm{Fr}_{\mathbf{SL}}(E,\tau)$ is a principal $\mathbf{SL}(r,\mathbb{C})$-bundle, whose fiber at $x\in X$ is given by:
@@
The right $\mathbf{SL}(r,\mathbb{C})$-action on the fiber of $x\in X$ is given by:
@@
Following the discussion after Lemma 2.3, stability for vector bundles $E$ and special vector bundles $(E,\tau)$ should be equivalent since Lemmas 2.3 and 2.6 give the analogous correspondences. The decoration of a reduction of $E$ to $\mathbf{SL}(r,\mathbb{C})$ does not restrict the set of subbundles $F\subsetneq E$ with respect to which one tests the slope conditions. This means that the stability and semistability conditions for decorated vector bundles $(E,\tau)$, where $\tau$ is a non-vanishing section of $\det(E)$, are exactly that of stability and semistability of the underlying vector bundle $E$. Observe that $\mathbf{SL}(r,\mathbb{C})$ is the commutator subgroup $\mathbf{GL}(r,\mathbb{C}),\mathbf{GL}(r,\mathbb{C})]$ of $\mathbf{GL}(r,\mathbb{C})$. In Lemma 2.13, we will see the same phenomenon as Lemma 2.6, that the stability of principal $[G,G]$-bundles and the stability of their extensions to principal $G$-bundles are equivalent. Note that the commutator subgroup $[G,G]$ is always semisimple and has the same root system as $G$.
@@ 2.2. Ramanathan-semistability
We review Ramanathan’s work in [Ram75], including his definition of semistability and stability for principal $G$-bundles, where $G$ is complex and reductive. Let $\xi$ be a holomorphic principal $G$-bundle on a compact Riemann surface $X$. For a reduction $s:X\rightarrow\xi/H$ of $\xi$ to a subgroup $H\subseteq G$, we can pull the bundle $\xi\rightarrow\xi/H$ back by $s$ to obtain a principal $H$-bundle $s^{*}\xi$ on $X$. We also refer to $s^{*}\xi$ as the reduction of $\xi$ to $H$. Following [Ram75], we define the vertical tangent bundle $V_{\xi/H}$, which is the vector bundle on $\xi/H$ consisting of tangent vectors along the fibers of $\xi/H\rightarrow X$.
@@
Let $\xi$ be a holomorphic principal $G$-bundle on a compact Riemann surface $X$.
@@
as seen in [Ram75, Remark 2.2]. Since $G$ is reductive, there exists a nondegenerate adjoint-invariant form on $\mathrm{Lie}(G)$ such that $\mathrm{ad}(\xi)$ is self-dual, so $\mathrm{ad}(\xi)$ has degree 0. From the short exact sequence, we have $\deg(s^{*}V_{\xi/P})=-\deg(\mathrm{ad}(s^{*}\xi))$. This can be used as an alternate characterization of Ramanathan-(semi)-stability. Namely, $\xi$ is Ramanathan-(semi)-stable when for all reductions $s:X\rightarrow\xi/P$ to maximal parabolic subgroups $P$ of $G$, we have $\deg(\mathrm{ad}(s^{*}\xi))(\leqslant)0$.
@@
Once we fix a Borel subgroup $B$, to verify Ramanathan-(semi)-stability, it suffices to verify the inequality for reductions to maximal standard parabolic subgroups of $G$. For a principal $G$-bundle $\xi$ with this property and a reduction $s:X\rightarrow\xi/P$ to any maximal parabolic subgroup $P\subsetneq G$, we have that:
@@
    There exists $g\in G$, such that $gPg^{-1}=P_{I}$ for some standard maximal parabolic subgroup $P_{I}\subsetneq G$. See [Bor91, IV. 11.1 Theorem, IV. 11.3 Corollary].
@@
The next lemma is from [HM04, Proposition 1], which we use to recall the equivalence between slope-(semi)-stability and Ramanathan-(semi)-stability. Let $E$ be a vector bundle of rank $r$.
@@
We can now prove the following theorem, from [HM04, Corollary 1].
@@
We can prove Theorem 2.12 in the same way as Theorem 2.11. Alternatively, we can invoke the following lemma, which generalizes the connection between Theorems 2.11 and 2.12 and shows why the notion of stability and semistability of a principal $[G,G]$-bundle coincides with that of the associated $G$-bundle.
@@
For a principal $[G,G]$-bundle $\xi_{1}$, the following are equivalent:
@@
Let $P_{1}$ be a maximal parabolic subgroup of $[G,G]$. Let $P=P_{1}\mathbf{R}(G)$ be the subgroup of $G$ generated by $P_{1}$ and the radical $\mathbf{R}(G)$ of $G$, as defined in [Bor91, IV. 11.21], then $P$ is a maximal parabolic subgroup of $G$, since $G=[G,G]\mathbf{R}(G)$ following [Bor91, IV. 14.2 Proposition], and the fact that $\mathbf{R}(G)$ is the connected component of the center $\mathbf{Z}(G)$ from [Bor91, IV. 11.21 Proposition]. Furthermore, we have a $[G,G]$-equivariant isomorphism of complete varieties:
@@
Lastly, we recall an important use of Remark 2.8. The adjoint bundle of a principal $G$-bundle determines its Ramanathan-semistability. To prove this, we use [RR84, Theorem 3.18].
@@
Let $\varphi:G\rightarrow H$ be a homomorphism of complex reductive groups that maps the radical $\mathbf{R}(G)$ into $\mathbf{R}(H)$. For a Ramanathan-semistable principal $G$-bundle $\xi$, the extension $\xi(H)$ of $\xi$ to $H$ is Ramanathan-semistable.
This theorem, proven in [RR84] by defining admissible reductions, will be helpful to us in later sections as well. We now present the result of [Ram96, 3.18 Corollary].
@@
A principal $G$-bundle $\xi$ is Ramanathan-semistable if and only if its adjoint bundle $\mathrm{ad}(\xi)$ is slope-semistable.
@@
The forward implication is a consequence of Theorem 2.14: the radical of a reductive group is a subgroup of the center, from [Bor91, IV. 11.21 Proposition], and therefore $\mathbf{R}(G)$ gets mapped to 0 under the adjoint representation, whose kernel is $\mathbf{Z}(G)$. Ramanathan-semistability is preserved in the frame bundle of the adjoint bundle $\mathrm{ad}(\xi)$, which is equivalent to slope-semistability due to Theorem 2.11.
@@
The reverse implication of Lemma 2.15 is also true for stability, as we can replace the inequalities with strict inequalities in the proof above. However, the same is not true for the forward implication. In [HM04, Lemma 2], it is explained that for stability, the forward implication is already false if $G$ is not semisimple, as in the case of $G=\mathbf{GL}(r,\mathbb{C})$.